# Relative Entropy Methods for Calculating Committors

### v1

@@ 3. A Change of Measure Formula for Approximations to the Transition Path Process
Since $\tilde{q}>0$ on $\mathscr{T}\cup\partial B$, $r$ is smooth in an open neighborhood of any point in $\mathscr{T}\cup\partial B$. It will suffice to show that $r$ is smooth up to $\partial A$. Let $x\in\mathscr{D}$. By Taylor’s theorem and Lemma 1,
@@
since the proof of Lemma 1 verifies that $\nabla\tilde{q}(\bar{x})=\lvert\nabla\tilde{q}(\bar{x})\rvert n$ under our assumptions on $\tilde{q}$. Therefore,
@@
The result follows since $\bar{x}$ and $h$ are smooth functions of $x$, $\tilde{q}$ and $q$ are in $C^{\infty}(\overline{\Omega})$, and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ are positive. ∎
Lemma 2 implies that $\log r$ and $\nabla\log r$ are smooth and bounded in a neighborhood of any point on $\partial A$, since $\lvert\nabla q(x)\rvert>0$ for all $x\in\partial A$ by Lemma 1. If $\nabla\log r$ is in fact globally bounded, then the Novikov condition stated in Assumption 5 below holds, and Girsanov’s theorem guarantees a change of measure.
Assumption 5.
@@
$\mathds{Q}\left[\exp\left(\varepsilon\int_{0}^{t}\lvert\nabla\log r(X_{s})\rvert^{2}\,\mathrm{d}s\right)\right]<\infty.$
Establishing or assuming a global bound on $\nabla\log r$ seems to us to be the most practical means of verifying the Novikov condition and the existence of a change of measure. For example, a global bound on $\nabla\log r$ follows from Assumption 4 and Lemma 2 when $\mathscr{T}$ is bounded. We note, however, that a global bound is only a convenient sufficient condition, not necessary, for the results below.
To state our change of measure, we require some additional notation to properly specify the probability space on which the exact and approximate transition path processes are defined. Let $B_{t}$ be a $d$-dimensional standard Brownian motion on the probability space $(\Omega,\mathcal{F},\mathds{Q})$, and suppose that $Y_{0}$ is a random variable on $(\Omega,\mathcal{F})$ taking values in $\overline{\mathscr{T}}$ and independent of $B_{t}$. Let $\mathcal{F}_{t}$ be the filtration generated by $B_{t}$ and $Y_{0}$. In [13], it is shown that a strong solution of
$dY_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$ (8)
@@
Lemma 3.
Let Assumptions 234, and 5 hold. Let $Y_{t}$ and $Y_{0}$ be as described in the previous paragraph. The process
$M_{t}:=\exp\left(\int_{0}^{t}\sqrt{2\varepsilon}\nabla\log r(Y_{s})\cdot\mathrm{d}B_{s}-\frac{1}{2}\int_{0}^{t}\lvert\sqrt{2\varepsilon}\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)$ (9)
is a positive $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{Q}[M_{t}]=1$ for all $t\geq 0$. Therefore, there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ so that for any fixed $T>0$,
@@
Proof.
The above merely restates Girsanov’s theorem. Observe that Girsanov’s theorem holds even when the drift is singular. Only some form of Novikov’s condition is required. ∎
Remark 3 (Weak Solutions of the Approximate Transition Path Equation).
Lemma 3 establishes the existence of at least a weak solution to
@@
00 0 $=-\nabla U(x)\cdot\nabla\log q(x)+\varepsilon\Delta\log q(x)+\varepsilon\lvert\nabla\log q(x)\rvert^{2}$
@@
on $x\in\mathbb{R}^{d}\setminus(A\cup B)$, which implies the result since $\exp(m)$ must be positive. ∎
We use the Hamilton–Jacobi–Bellman equation in Lemma 4 to eliminate $q$ from the exponential factor in formula (9) for $M_{t}$.
Lemma 5.
Under the hypotheses of Lemma 3, and assuming $Y_{0}$ takes values in $\partial A$, we have
@@
$=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\nabla\tilde{q}(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)$ (10)
@@
Applying the above with $g=\log r$ and using Lemma 2, we have for $t>0$ that
@@
×$\quad\times\exp\left(-\int_{0}^{t}L\log r(Y_{s})+2\varepsilon\nabla\log q(Y_{s})^{t}\nabla\log r(Y_{s})+\varepsilon\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right).$
By Lemma 4,
@@
For Theorem 1 below to hold, $\tau$ must be finite for both the exact and approximate transition path processes. Assumption 1 on the ergodicity of the overdamped Langevin dyanmics implies that the exact transition path process hits $B$, since $B$ has positive Lebesgue measure. That is, $\mathds{Q}[\tau<\infty]=1$. We assume that that $\tau$ is finite for the approximate process as well.
Assumption 6.
@@
for any initial value $Y_{0}$ supported in $\overline{\mathscr{T}}$. Recall that $\mathds{P}$ is the measure introduced in Lemma 3 under which $Y_{t}$ is a weak solution of the approximate transition path equation (7).
We eliminate $\lvert\nabla q(Y_{0})\rvert$ from (10) by taking $Y_{0}$ to have the reactive flux distribution under $\mathds{Q}$. In that case, the factor $\lvert\nabla q(Y_{0})\rvert$ in the reactive flux (5) cancels with the one in (10). Everything that remains in the change of measure can be computed without knowing the exact committor $q$, except for a normalizing constant. The result is formula (11) in Theorem 1.
Theorem 1.
@@
$Y_{0}\sim\frac{1}{\nu}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
@@
$Z_{t}:=\frac{\nu}{\mu}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right),$
@@
$Z_{0}:=\frac{\nu}{\mu}\frac{m(Y_{0})}{\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}.$
The process $Z_{t}$ is a nonnegative $\mathcal{F}_{t}$-martingale with $\mathds{Q}[Z_{t}]=1$ for all $t\geq 0$, so there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ with
@@
is a Brownian motion and $Y_{0}\sim\mu^{-1}m(x)\,\mathrm{d}S_{A}(x)$. The density of $\mathds{P}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$ is
$\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{\tau}}=Z_{\tau}=\frac{\nu}{\mu}\tilde{q}(Y_{\tau})\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right).$ (11)
Proof.
See Appendix A. ∎
4. Applications of the Change of Measure
In this section, we consider two applications of the change of measure formula: selection and training. By training, we mean the minimization of the relative entropy $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ of the approximate transition path process with respect to the exact process over a family of approximate committor functions. By selection, we mean the comparison of different approximations to the committor function based on relative entropy differences. This is somewhat similar to model selection based on the Bayes and Akaike information criteria except that we do not consider penalizing model complexity. We also speculate on two other possibilities: error estimation and importance sampling. By error estimation, we mean the direct estimation of $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$. By importance sampling, we mean the unbiased estimation of observables of the exact process given a sample of paths of the approximate process. Error estimation and importance sampling entail difficulties that we cannot completely resolve at this time; see Section 5 for details.
4.1. Relative Entropy
@@
Definition 2 (Relative entropy).
@@
$D_{\mathrm{KL}}(P\|Q):=Q\left[\frac{\mathrm{d}P}{\mathrm{d}Q}\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right]=P\left[\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right].$
@@
By abuse of notation, throughout the rest of this work, we will write $\mathds{P}$ and $\mathds{Q}$ for the restrictions of $\mathds{P}$ and $\mathds{Q}$ to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. In other words, we will stop observing the process $Y_{t}$ when it hits $B$ at time $\tau$. By formula (11) in Theorem 1, we have
$D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ $=\mathds{P}\left[\log\left(\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right)\right]$
$=\log\nu-\log\mu$
$\qquad+\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\mu}\,\mathrm{d}S_{A}(x)$
$\qquad+\mathds{P}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$ (12)
@@
$D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$ $=\mathds{Q}\left[\log\left(\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}}\right)\right]$
$=\log\mu-\log\nu$
$\qquad+\int_{\partial A}\log\left(\frac{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}{m(x)}\right)\frac{\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)}{\nu}\,\mathrm{d}S_{A}(x)$
$\qquad+\mathds{Q}\left[\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s-\log\tilde{q}(Y_{\tau})\right].$ (13)
@@
Remark 4.
The existence of the change of measure $\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}$ does not imply $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$. We note that $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$ under the same conditions that imply finite variance of the estimator in Lemma 8, but with $\mathds{P}[\tau^{2}]<\infty$ replaced by $\mathds{P}[\tau]<\infty$. We leave the proof to the reader.
4.2. Selection: Estimating Relative Entropy Differences

### v2

@@ 3 A Change of Measure Formula for Approximations to the Transition Path Process
Since $\tilde{q}>0$ on $\mathscr{T}\cup\partial B$, $r$ is smooth in an open neighborhood of any point in $\mathscr{T}\cup\partial B$. It will suffice to show that $r$ is smooth up to $\partial A$. Let $x\in\mathscr{D}$. By Taylor’s theorem and Lemma 2.2,
@@
since the proof of Lemma 2.2 verifies that $\nabla\tilde{q}(\bar{x})=\lvert\nabla\tilde{q}(\bar{x})\rvert n$ under our assumptions on $\tilde{q}$. Therefore,
@@
The result follows since $\bar{x}$ and $h$ are smooth functions of $x$, $\tilde{q}$ and $q$ are in $C^{\infty}(\overline{\mathscr{T}})$, and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ and $\lvert\nabla\tilde{q}(\bar{x})\rvert$ are positive.
Lemma 3.1 implies that $\log r$ and $\nabla\log r$ are smooth and bounded in a neighborhood of any point on $\partial A$, since $\lvert\nabla q(x)\rvert>0$ for all $x\in\partial A$ by Lemma 2.2. If $\nabla\log r$ is in fact globally bounded, then the Novikov condition stated in Assumption 3 below holds, and Girsanov’s theorem guarantees a change of measure.
{assumption}
@@
$\mathds{E}_{\mathds{Q}}\left[\exp\left(\varepsilon\int_{0}^{t}\lvert\nabla\log r(X_{s})\rvert^{2}\,\mathrm{d}s\right)\right]<\infty.$
Establishing or assuming a global bound on $\nabla\log r$ seems to us to be the most practical means of verifying the Novikov condition and the existence of a change of measure. For example, a global bound on $\nabla\log r$ follows from Assumption 3 and Lemma 3.1 when $\mathscr{T}$ is bounded. We note, however, that a global bound is only a convenient sufficient condition, not necessary, for the results below.
To state our change of measure, we require some additional notation to properly specify the probability space on which the exact and approximate transition path processes are defined. Let $B_{t}$ be a $d$-dimensional standard Brownian motion on the probability space $(\Omega,\mathcal{F},\mathds{Q})$, and suppose that $Y_{0}$ is a random variable on $(\Omega,\mathcal{F})$ taking values in $\overline{\mathscr{T}}$ and independent of $B_{t}$. For now, we do not make any assumptions regarding the distribution of $Y_{0}$, but, in Theorem 3.11 below, $Y_{0}$ will have the reactive flux distribution (6). Let $\mathcal{F}_{t}$ be the filtration generated by $B_{t}$ and $Y_{0}$. In [24], it is shown that a strong solution of
(9) $dY_{t}=-\nabla U(Y_{t})\,\mathrm{d}t+2\varepsilon\nabla\log q(Y_{t})\,\mathrm{d}t+\sqrt{2\varepsilon}\,\mathrm{d}B_{t}$
@@
Lemma 3.3.
Let Assumptions 223, and 3 hold. Let $Y_{t}$ and $Y_{0}$ be as described in the previous paragraph. The process
(10) $M_{t}:=\exp\left(\int_{0}^{t}\sqrt{2\varepsilon}\nabla\log r(Y_{s})\cdot\mathrm{d}B_{s}-\frac{1}{2}\int_{0}^{t}\lvert\sqrt{2\varepsilon}\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right)$
is a positive $\mathcal{F}_{t}$-martingale under $\mathds{Q}$ with $\mathds{E}_{\mathds{Q}}[M_{t}]=1$ for all $t\geq 0$. Therefore, there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ so that for any fixed $T>0$,
@@
Proof 3.4.
The above merely restates Girsanov’s theorem; cf.[25, Section 8.6]. Observe that Girsanov’s theorem holds even when the drift is singular. Only some form of Novikov’s condition is required.
Remark 3.5 (Weak Solutions of the Approximate Transition Path Equation).
Lemma 3.3 establishes the existence of at least a weak solution to
@@
$0$ $=-\nabla U(x)\cdot\nabla\log q(x)+\varepsilon\Delta\log q(x)+\varepsilon\lvert\nabla\log q(x)\rvert^{2}$
@@
on $x\in\mathbb{R}^{d}\setminus(A\cup B)$, which implies the result since $\exp(m)$ must be positive.
We use the Hamilton–Jacobi–Bellman equation in Lemma 3.6 to eliminate $q$ from the exponential factor in formula (10) for $M_{t}$.
Lemma 3.8.
Under the hypotheses of Lemma 3.3, and assuming $Y_{0}$ takes values in $\partial A$, we have
@@
(11) $=\frac{\lvert\nabla q(Y_{0})\rvert}{\lvert\nabla\tilde{q}(Y_{0})\rvert}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right)$
@@
Applying the above with $g=\log r$ and using Lemma 3.1, we have for $t>0$ that
@@
$\quad\times\exp\left(-\int_{0}^{t}L\log r(Y_{s})+2\varepsilon\nabla\log q(Y_{s})^{t}\nabla\log r(Y_{s})+\varepsilon\lvert\nabla\log r(Y_{s})\rvert^{2}\,\mathrm{d}s\right).$
By Lemma 3.6,
@@
For Theorem 3.11 below to hold, $\tau$ must be finite for both the exact and approximate transition path processes. Assumption 2 on the ergodicity of the overdamped Langevin dynamics implies that the exact transition path process will hit $B$ with probability one, since $B$ has positive Lebesgue measure. That is, $\mathds{Q}[\tau<\infty]=1$. We assume that $\tau$ is finite for the approximate process as well.
{assumption}
@@
for any initial value $Y_{0}$ supported in $\overline{\mathscr{T}}$. Recall that $\mathds{P}$ is the measure introduced in Lemma 3.3 under which $Y_{t}$ is a weak solution of the approximate transition path equation (8).
We eliminate $\lvert\nabla q(Y_{0})\rvert$ from (11) by taking $Y_{0}$ to have the reactive flux distribution under $\mathds{Q}$. In that case, the factor $\lvert\nabla q(Y_{0})\rvert$ in the reactive flux (6) cancels with the one in (11). Everything that remains in the change of measure can be computed without knowing the exact committor $q$, except for a normalizing constant. The result is formula (12) in Theorem 3.11.
Theorem 3.11.
@@
$Y_{0}\sim\frac{1}{\zeta}\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)\,\mathrm{d}S_{A}(x),$
@@
$Z_{t}:=\frac{\zeta}{\eta}\frac{\tilde{q}(Y_{t})}{q(Y_{t})}\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{t}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right),$
@@
$Z_{0}:=\frac{\zeta}{\eta}\frac{m(Y_{0})}{\lvert\nabla q(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}.$
The process $Z_{t}$ is a nonnegative $\mathcal{F}_{t}$-martingale with $\mathds{E}_{\mathds{Q}}[Z_{t}]=1$ for all $t\geq 0$, so there exists a probability measure $\mathds{P}$ on $(\Omega,\mathcal{F})$ with
@@
is a Brownian motion and $Y_{0}\sim\eta^{-1}m(x)\,\mathrm{d}S_{A}(x)$. The density of $\mathds{P}$ restricted to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$ is
(12) $\left.\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right\rvert_{\mathcal{F}_{\tau}}=Z_{\tau}=\frac{\zeta}{\eta}\tilde{q}(Y_{\tau})\frac{m(Y_{0})}{\lvert\nabla\tilde{q}(Y_{0})\rvert\exp(-U(Y_{0})/\varepsilon)}\exp\left(-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right).$
Proof 3.12.
See Appendix A.
4 Applications of the Change of Measure
In this section, we consider two applications of the change of measure formula: selection and training. By training, we mean the minimization of the relative entropy $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ of the approximate transition path process with respect to the exact process over a family of approximate committor functions. By selection, we mean the comparison of different approximations to the committor function based on relative entropy differences. This is somewhat similar to model selection based on the Bayes and Akaike information criteria except that we do not consider penalizing model complexity. We also speculate on two other possibilities: error estimation and importance sampling. By error estimation, we mean the direct estimation of $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ given a sample of trajectories from the approximate transition path process. By importance sampling, we mean the unbiased estimation of observables of the exact process given a sample of paths of the approximate process. Error estimation and importance sampling entail theoretical difficulties that we cannot resolve at this time, but see Section 6 for a simple example where both strategies appear to work well.
4.1 Relative Entropy
@@
Definition 4.1 (Relative entropy).
@@
$D_{\mathrm{KL}}(P\|Q):=\mathds{E}_{Q}\left[\frac{\mathrm{d}P}{\mathrm{d}Q}\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right]=\mathds{E}_{P}\left[\log\left(\frac{\mathrm{d}P}{\mathrm{d}Q}\right)\right].$
@@
By abuse of notation, throughout the rest of this work, we will write $\mathds{P}$ and $\mathds{Q}$ for the restrictions of $\mathds{P}$ and $\mathds{Q}$ to the stopping time $\sigma$-algebra $\mathcal{F}_{\tau}$. In other words, we will stop observing the process $Y_{t}$ when it hits $B$ at time $\tau$. By formula (12) in Theorem 3.11, we have
$D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ $=\mathds{E}_{\mathds{P}}\left[\log\left(\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}\right)\right]$
$=\log\zeta-\log\eta$
$\qquad+\int_{\partial A}\log\left(\frac{m(x)}{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}\right)\frac{m(x)}{\eta}\,\mathrm{d}S_{A}(x)$
(13) $\qquad+\mathds{E}_{\mathds{P}}\left[\log\tilde{q}(Y_{\tau})-\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s\right]$
@@
$D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$ $=\mathds{E}_{\mathds{Q}}\left[\log\left(\frac{\mathrm{d}\mathds{Q}}{\mathrm{d}\mathds{P}}\right)\right]$
$=\log\eta-\log\zeta$
$\qquad+\int_{\partial A}\log\left(\frac{\lvert\nabla\tilde{q}(x)\rvert\exp(-U(x)/\varepsilon)}{m(x)}\right)\frac{\lvert\nabla q(x)\rvert\exp(-U(x)/\varepsilon)}{\zeta}\,\mathrm{d}S_{A}(x)$
(14) $\qquad+\mathds{E}_{\mathds{Q}}\left[\int_{0}^{\tau}\frac{L\tilde{q}}{\tilde{q}}(Y_{s})\,\mathrm{d}s-\log\tilde{q}(Y_{\tau})\right].$
@@
Remark 4.2 (Which Entropy?).
In the present work, we focus on $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ instead of $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$, since $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})$ can be estimated or minimized given a sample of approximate reactive trajectories drawn from $\mathds{P}$. If one instead had a sample of exact reactive trajectories from $\mathds{Q}$, then one would prefer $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P})$. A sample of exact reactive trajectories could possibly come from a very long molecular dynamics simulation of a not very metastable system or from an enhanced sampling method like parallel replica dynamics [32]. We leave it as an exercise for the reader to show that if $q_{\theta}$ is a smooth parametric family of functions that satisfy Assumption 3 and $\mathds{P}_{\theta}$ is the distribution of the transition path process with $q_{\theta}$ in place of $q$, then $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$ is differentiable under much weaker conditions on $q_{\theta}$ than Assumption 4.3, which implies differentiability of $D_{\mathrm{KL}}(\mathds{P}_{\theta}\|\mathds{Q})$. Moreover, one can easily estimate $\nabla_{\theta}D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$ given a sample from $\mathds{Q}$. Therefore, given a sample from $\mathds{Q}$ and knowledge of the generator $L$, one could train an approximate committor to minimize $D_{\mathrm{KL}}(\mathds{Q}\|\mathds{P}_{\theta})$.
Remark 4.3.
The existence of the change of measure $\frac{\mathrm{d}\mathds{P}}{\mathrm{d}\mathds{Q}}$ does not imply $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$. We note that $D_{\mathrm{KL}}(\mathds{P}\|\mathds{Q})<\infty$ under the same conditions that imply finite variance of the estimator in Lemma B.2, but with $\mathds{E}_{\mathds{P}}[\tau^{2}]<\infty$ replaced by $\mathds{E}_{\mathds{P}}[\tau]<\infty$. We leave the proof to the reader.
4.2 Selection: Estimating Relative Entropy Differences